If the production possibilities frontiers shown are each for two days of production, then which of the following combinations of bolts and nails could Uzbekistan and Azerbaijan together not make in a given 2-day production period?
Answer #1

Answer

The correct answer is (d) 47 bolts and 18 nails

We can see from above production function that both functions opportunity cost of producing 30 bolts for Uzbekistan is 60 nails.

Hence, opportunity cost of producing 1 bolts for Uzbekistan is 60/30 = 2 nails. Similarly opportunity cost of producing 1 bolts for Azerbaijan is 80/20 = 4 nails.

Hence As opportunity cost of producing 1 bolts in Uzbekistan < opportunity cost of producing 1 bolt in Azerbaijan.

Hence Uzbekistan will specialize in Bolts and Azerbaijan will specialize in nails.

Now lets check which option is incorrect:

(a) 9 bolts and 122 nails - As Azerbaijan will specialize in nails, It will use its entire time on Nails and will produce 80 nails. We can see from the PPF that Uzbekistan can produce 9 bolts and 42 nails. Hence Option (a) is achievable

(b) 21 bolts and 98 nails - As Azerbaijan will specialize in nails, It will use its entire time on Nails and will produce 80 nails. We can see from the PPF that Uzbekistan can produce 21 bolts and 18 nails. Hence Option (b) is achievable

(c) 36 bolts and 56 nails - Uzbekistan will specialize in Bolts, It will use its entire time on bolts and will produce 30 bolts. We can see from the PPF that Azerbaijan can produce 6 bolts and 56 nails. Hence Option (c) is achievable

(d) 47 bolts and 18 nails - Uzbekistan will specialize in Bolts, It will use its entire time on bolts and will produce 30 bolts. We can see from the PPF that Azerbaijan can produce 17 bolts and 12 nails. Hence Option (d) is not achievable. Hence option (d) is incorect.

Note :

PPF equation of Uzbekistan is N + 2B = 60

PPF equation of Azerbaijan is N + 4B = 80